-3b+ac=c-4
Solve for c

Answer by persian52(161)

-3b+ac=c-4
Since c contains the variable to solve for, move it to the left-hand side of the equation by subtracting c from both sides.
ac-3b-c=-4
Move all terms not containing c to the right-hand side of the equation.
ac-c-3b=-4
Factor out the GCF of c from each term in the polynomial.
c(a)+c(-1)=3b-4
Factor out the GCF of c from ac-c.
c(a-1)=3b-4
Divide each term in the equation by (a-1).
(c(a-1))/(a-1)=(3b)/(a-1)-(4)/(a-1)
Simplify the left-hand side of the equation by canceling the common terms.
c=(3b)/(a-1)-(4)/(a-1)
Simplify the right-hand side of the equation by simplifying each term.
c=(3b-4)/(a-1)